# MiLoRA: Harnessing Minor Singular Components for Parameter-Efficient LLM Finetuning

### v2

@@ 5.2 Is Minor Singular Component Adaptation Important?
To investigate the influence of singular components of varying magnitudes on finetuning performance, we conduct experiments on math reasoning tasks using the LLaMA2-7B model. Specifically, we finetune LLaMA2-7B on the MetaMathQA 395K dataset for 1 epoch with a maximum sequence length of 512 to save computation. We initialize the low-rank matrices ${\bm{A}}$ and ${\bm{B}}$ with principal, minor, and randomly sampled singular components, and report the evaluation results in Table 6. As can be seen, initializing with the minor singular components achieves the best performance across both datasets. This underscores the importance of adapting the minor singular components during the finetuning process. Concurrent with our work, Meng et al. (2024) propose PiSSA, a similar approach to MiLoRA but adapts the principal singular components. We will provide a detailed comparison between MiLoRA and PiSSA in Section 5.5.
@@
5.3 How does the Matrix Update $\Delta W$ Compare to $W$?
We investigate the relationship between $\Delta{\bm{W}}$ and ${\bm{W}}$ of different methods, where $\Delta{\bm{W}}$ is the difference between the finetuned and pre-trained weights. Following the analysis method in Hu et al. (2021), we project ${\bm{W}}$ and $\Delta{\bm{W}}$ onto the $r$-dimensional subspace of $\Delta{\bm{W}}$ by computing ${\bm{U}}^{T}{\bm{W}}{\bm{V}}$ and ${\bm{U}}^{T}\Delta{\bm{W}}{\bm{V}}$, with ${\bm{U}}$/${\bm{V}}$ being the left/right top $r$ singular-vector matrix of $\Delta{\bm{W}}$. As a comparison, we also compute ${\bm{U}}^{T}{\bm{W}}{\bm{V}}$ by replacing ${\bm{U}}$, ${\bm{V}}$ with the top $r$ singular vectors of ${\bm{W}}$ or a random matrix. We use the Frobenius norm as the measurement of magnitude. The results are shown in Table 7. Following Hu et al. (2021), we use the query weight in the middle layer (16th) of the model for analysis. We make two key observations. First, $\Delta{\bm{W}}$ of LoRA and MiLoRA have a similar correlation with ${\bm{W}}$, both slightly stronger than a random matrix, indicating that they amplify directions not emphasized in ${\bm{W}}$. In contrast, $\Delta{\bm{W}}$ of PiSSA shows a strong correlation with ${\bm{W}}$. This is attributed to PiSSA’s direct optimization of the principal singular components. Second, the amplification factor varies significantly across different methods. Compared to LoRA, we find the amplification factor of MiLoRA is significantly smaller. This indicates that MiLoRA has a reduced impact on the features which already present in ${\bm{W}}$. See Table 13 of Appendix B for analyzing the weight matrix of MLP down projection in the same layer, which demonstrates similar trends.
@@ 5.3 How does the Matrix Update $\Delta W$ Compare to $W$?
To further analyze the characteristics of $\Delta{\bm{W}}$ for different methods, we measure the normalized subspace similarity (Hu et al., 2021) between the subspaces spanned by top $r$ singular vectors of $\Delta{\bm{W}}$ and the subspace spanned by top $r$ singular vectors, bottom $r$ singular vectors, and a random $r$ singular vectors of ${\bm{W}}$. We define the normalized similarity metric based on the Grassmann distance following Hu et al. (2021) (See Appendix A.1 for more details). The results are presented in Figure 2. From the left figure, it is evident that PiSSA focuses on optimizing features associated with the top singular vectors, whereas LoRA and MiLoRA do not significantly optimize these features. MiLoRA primarily optimizes features related to the bottom singular vectors. The right figures further show that MiLoRA places greater emphasis on features associated with the bottom singular vectors of the MLP layers.
@@ 5.5 Comparison Between MiLoRA and PiSSA
Motivation: PiSSA is designed to approximate full finetuning by adapting the principal singular components, which are believed to capture the essence of the weight matrices. In contrast, our method MiLoRA aims to adapt to new tasks while maximally retaining the base model’s knowledge. To achieve this, we instead finetune the minor singular components of the weight matrices, which are less important for the weight matrices.
Performance: The PiSSA paper claims that PiSSA outperforms both LoRA and adapting minor singular components. However, our investigation suggests that this claim is likely due to the specific hyperparameters used in their experiments. To verify this, we replicate their experimental setup using the same MetaMathQA 100K dataset and compare the performance of PiSSA, LoRA, and MiLoRA under both PiSSA’s hyperparameters and our own. Specifically, we choose a rank of 128 where PiSSA obtains the best performance in their paper and finetunes for 1 epoch for all setups following PiSSA paper. The results are summarized in Table 9. We have several observations. First, under the PiSSA hyperparameter setup, PiSSA does indeed outperform both LoRA and MiLoRA, consistent with the claims made in the original PiSSA paper. Second, we find that all methods perform better in our hyperparameter setup, suggesting that our hyperparameter configuration is more effective. PiSSA configuration of hyperparameters has a combination of a small learning rate (2e-5) and a large batch size (128). This could potentially result in slow learning speed and sub-optimal performance for common PEFTs, as the optimal learning rate for common PEFTs is generally much higher than that for full finetuning (Lialin et al., 2023; Biderman et al., 2024). Third, in our hyperparameter setup, MiLoRA performs the best, outperforming LoRA and PiSSA by 1.4 and 4.3 average scores, respectively. In summary, when using an appropriate hyperparameter configuration, our proposed MiLoRA method achieves markedly superior performance compared to PiSSA.
@@ 6 Related Work
After pretraining, LLMs are finetuned with instruction data or human preference data to adapt to different downstream tasks (Ouyang et al., 2022; Yu et al., 2023; Cui et al., 2023). Parameter-efficient finetuning (Lialin et al., 2023, PEFT) explores effective approaches to reduce the computational resources required during finetuning while managing to match the performance of full finetuning. Previous PEFT methods can be grouped into three lines: adapter-, LoRA- and prompt-based methods. Adapter-based methods insert additional trainable feedforward submodules into each Transformer layer (Houlsby et al., 2019; Pfeiffer et al., 2021a, b). However, they introduce additional computational costs during inference. LoRA-based methods (Hu et al., 2021; Liu et al., 2024b; Zhang et al., 2023) model the changes of selected linear weights as low-rank matrices. During finetuning, given a linear matrix, only two low-rank matrices are optimized while the pretrained weight matrix is frozen. Prompt-based methods (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2022) add additional soft prompts to the input tokens. During training, only soft prompts are updated while the pretrained model parameters are fixed. Among these PEFT methods, the LoRA-based approaches are widely used for LLM fine-tuning because they are easy to implement and do not introduce computational overhead during inference.
@@
LoRA (Hu et al., 2021) reparameterizes the weight update with two trainable low-rank matrices, while freezing the pretrained weights. With this lightweight decomposition, LoRA reduces storage and task-switching overhead by sharing the pretrained models across multiple tasks. Since then, researchers have explored and proposed different LoRA variants for PEFT. AdaLoRA (Zhang et al., 2023) and ALoRA (Liu et al., 2024d), adaptively determine the rank of LoRA module in each weight matrix according to the importance score. The rsLoRA (Kalajdzievski, 2023) modifies the LoRA with the appropriate scaling factor to improve the performance of large ranks. The DoRA (Liu et al., 2024b) method decomposes the pretrained weight into the magnitude and directional components, then finetunes both for better performance. Concurrent with our work, PISSA (Meng et al., 2024) and LoRA-GA Wang et al. (2024b) propose to better approximate full finetuning by only updating the principal singular components or aligning the gradients of low-rank updates to that of full finetuning.
@@ 7 Conclusion
Due to limited computational resources, we only examine the effectiveness of MiLoRA on the LLaMA-family LLMs and tasks like commonsense reasoning, math reasoning and instruction following benchmarks following previous work Liu et al. (2024b); Wu et al. (2024b). We examine and compare MiLoRA with baselines using the hyperparameter configurations provided in the previous works Hu et al. (2023); Liu et al. (2024b) instead of exhaustive hyperparameter search for each task. We leave the exploration of MiLoRA on other tasks and other LLMs like Mistral and Gemma as future work.
@@ A.3 PiSSA Hyperparameter Setup for LLM
A.4 More Implementation Details
We use the implementation of LLM-Adapters (Hu et al., 2023) [^2] https://github.com/AGI-Edgerunners/LLM-Adapters for commonsense reasoning, the implementation of PiSSA (Meng et al., 2024) [^3] https://github.com/GraphPKU/PiSSA for math reasoning and the implementation of open-instruct [^4] https://github.com/allenai/open-instruct for instruction tuning. For hyperparameters of finetuning the vision-language model, we follow the LoRA configuration provided by Liu et al. (2024a) without tuning for all PEFT methods for a fair comparison. Table 11 shows the details.

### v3

@@ 5.2 Is Minor Singular Component Adaptation Important?
To investigate the influence of singular components of varying magnitudes on finetuning performance, we conduct experiments on math reasoning tasks using the LLaMA2-7B model. Specifically, we finetune LLaMA2-7B on the MetaMathQA 395K dataset for 1 epoch with a maximum sequence length of 512 to save computation. We initialize the low-rank matrices ${\bm{A}}$ and ${\bm{B}}$ with principal, minor, and randomly sampled singular components, and report the evaluation results in Table 6. As can be seen, initializing with the minor singular components achieves the best performance across both datasets. This underscores the importance of adapting the minor singular components during the finetuning process.
@@
5.3 How does the Matrix Update $\Delta W$ Compare to $W$?
We investigate the relationship between $\Delta{\bm{W}}$ and ${\bm{W}}$ of different methods, where $\Delta{\bm{W}}$ is the difference between the finetuned and pre-trained weights. Following the analysis method in Hu et al. (2021), we project ${\bm{W}}$ and $\Delta{\bm{W}}$ onto the $r$-dimensional subspace of $\Delta{\bm{W}}$ by computing ${\bm{U}}^{T}{\bm{W}}{\bm{V}}$ and ${\bm{U}}^{T}\Delta{\bm{W}}{\bm{V}}$, with ${\bm{U}}$/${\bm{V}}$ being the left/right top $r$ singular-vector matrix of $\Delta{\bm{W}}$. As a comparison, we also compute ${\bm{U}}^{T}{\bm{W}}{\bm{V}}$ by replacing ${\bm{U}}$, ${\bm{V}}$ with the top $r$ singular vectors of ${\bm{W}}$ or a random matrix. The results are shown in Table 7. Following Hu et al. (2021), we use the query weight in the middle layer (16th) of the model for analysis and use the Frobenius norm as the measurement of magnitude.
We make two key observations. First, $\Delta{\bm{W}}$ of LoRA and MiLoRA have a similar correlation with ${\bm{W}}$, both slightly stronger than a random matrix, indicating that they amplify directions not emphasized in ${\bm{W}}$. In contrast, $\Delta{\bm{W}}$ of PiSSA shows a strong correlation with ${\bm{W}}$. This is attributed to PiSSA’s direct optimization of the principal singular components. Second, the amplification factor varies significantly across different methods. Compared to LoRA, we find the amplification factor of MiLoRA is significantly smaller. This indicates that MiLoRA has a reduced impact on the features which already present in ${\bm{W}}$. See Table 14 of Appendix B.2 for analyzing the weight matrix of MLP down projection in the same layer, which demonstrates similar trends.
To further analyze the characteristics of $\Delta{\bm{W}}$ for different methods, we measure the normalized subspace similarity (Hu et al., 2021) between the subspaces spanned by top $r$ singular vectors of $\Delta{\bm{W}}$ and the subspace spanned by top $r$ singular vectors, bottom $r$ singular vectors, and a random $r$ singular vectors of ${\bm{W}}$. We define the normalized similarity metric based on the Grassmann distance following Hu et al. (2021) (See Appendix A.1 for more details) and report the results in Figure 2. From the left subfigure, it is clear that PiSSA learns in directions closely aligned with the top singular vectors of the pretrained weights, while LoRA and MiLoRA do not significantly optimize in these directions. Specifically, MiLoRA focuses on optimizing directions associated with the bottom singular vectors of the pretrained weights. The right subfigure further shows that weight matrices in both MLP and self attention modules of MiLoRA learn in directions more closely aligned with the bottom singular vectors, especially the MLP modules.
@@ 5.5 Comparison Between MiLoRA and PiSSA
Motivation: PiSSA is designed to approximate full finetuning by adapting the principal singular components, which are believed to capture the essence of the weight matrices. In contrast, our method MiLoRA aims to adapt to new tasks while maximally retaining the base model’s knowledge. To achieve this, we instead finetune the minor singular components of the weight matrices, which are less important for the pretrained weight matrices.
Performance: The PiSSA paper claims that PiSSA outperforms both LoRA and adapting minor singular components. However, our investigation suggests that this claim is likely due to the specific hyperparameters used in their experiments. To verify this, we replicate their experimental setup using the same MetaMathQA 100K dataset and compare the performance of PiSSA, LoRA, and MiLoRA under both PiSSA’s hyperparameters and our own. Specifically, we choose a rank of 128 where PiSSA obtains the best performance in their paper and finetunes for 1 epoch for all setups following PiSSA paper. The results are summarized in Table 9.
We have several observations. First, under the PiSSA hyperparameter setup, PiSSA does indeed outperform both LoRA and MiLoRA, consistent with the claims made in the original PiSSA paper. Second, we find that all methods perform better in our hyperparameter setup, suggesting that our hyperparameter configuration is more effective. PiSSA configuration of hyperparameters has a combination of a small learning rate (2e-5) and a large batch size (128). This could potentially result in slow learning speed and sub-optimal performance for common PEFTs, as the optimal learning rate for common PEFTs is generally much higher than that for full finetuning (Lialin et al., 2023; Biderman et al., 2024). Third, in our hyperparameter setup, MiLoRA performs the best, outperforming LoRA and PiSSA by 1.4 and 4.3 average scores, respectively. In summary, when using an appropriate hyperparameter configuration, our proposed MiLoRA method achieves markedly superior performance compared to PiSSA.
@@ 6 Related Work
After pretraining, LLMs are finetuned with instruction data or human preference data to adapt to different downstream tasks (Ouyang et al., 2022; Yu et al., 2023; Cui et al., 2023). Parameter-efficient finetuning (Lialin et al., 2023, PEFT) explores effective approaches to reduce the computational resources required during finetuning while managing to match the performance of full finetuning. Previous PEFT methods can be grouped into three lines: adapter-, LoRA- and prompt-based methods. Adapter-based methods insert additional trainable feedforward submodules into each Transformer layer (Houlsby et al., 2019; Pfeiffer et al., 2021a, b). However, they introduce additional computational costs during inference. LoRA-based methods (Hu et al., 2021; Liu et al., 2024b; Zhang et al., 2023; Wang et al., 2024a) model the changes of selected linear weights as low-rank matrices. During finetuning, given a linear matrix, only two low-rank matrices are optimized while the pretrained weight matrix is frozen. Prompt-based methods (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2022) add additional soft prompts to the input tokens. During training, only soft prompts are updated while the pretrained model parameters are fixed. Among these PEFT methods, the LoRA-based approaches are widely used for LLM fine-tuning because they are easy to implement and do not introduce computational overhead during inference.
@@
LoRA (Hu et al., 2021) reparameterizes the weight update with two trainable low-rank matrices, while freezing the pretrained weights. With this lightweight decomposition, LoRA reduces storage and task-switching overhead by sharing the pretrained models across multiple tasks. Since then, researchers have explored and proposed different LoRA variants for PEFT. AdaLoRA (Zhang et al., 2023) and ALoRA (Liu et al., 2024d), adaptively determine the rank of LoRA module in each weight matrix according to the importance score. The rsLoRA (Kalajdzievski, 2023) modifies the LoRA with the appropriate scaling factor to improve the performance of large ranks. The DoRA (Liu et al., 2024b) method decomposes the pretrained weight into the magnitude and directional components, then finetunes both for better performance. Concurrent with our work, PISSA (Meng et al., 2024) and LoRA-GA Wang et al. (2024c) propose to better approximate full finetuning by only updating the principal singular components or aligning the gradients of low-rank updates to that of full finetuning.
@@ 7 Conclusion
Due to limited computational resources, we primarily evaluate the effectiveness of MiLoRA on LLaMA-family LLMs and LLaVA-1.5, focusing on tasks such as commonsense reasoning, math reasoning, and instruction following, in line with prior work Liu et al. (2024b); Meng et al. (2024); Wu et al. (2024b). We examine and compare MiLoRA with baselines using the hyperparameter configurations provided in the previous works Hu et al. (2023); Liu et al. (2024a) instead of an exhaustive hyperparameter search for each task. We leave the exploration of MiLoRA on other tasks and other LLMs like Mistral and Gemma as future work.
Acknowledgements
This project was supported by National Natural Science Foundation of China (No. 62306132, No. 62106138) and Guangdong Basic and Applied Basic Research Foundation. We thank the anonymous reviewers for their insightful feedback on this work.
@@ A.3 PiSSA Hyperparameter Setup for LLM
A.4 Hyperparameter Setup for Vision-Language Model
For hyperparameters of finetuning the vision-language model, we follow the LoRA configuration provided by Liu et al. (2024a) without tuning for all PEFT methods for a fair comparison. Table 11 shows the details.
@@ Appendix B More Experiment Results
B.1 Experiments on Qwen2.5-7B
Method GSM8K MATH Avg.
LoRA 81.8 49.3 65.5
PiSSA 81.7 45.4 63.6
MiLoRA 85.5 48.7 67.1
Table 13: Math reasoning evaluation results for Qwen2.5-7B on GSM8K and MATH.
To enhance the credibility of our experiments, we evaluate MiLoRA using Qwen2.5-7B on the MetaMathQA-100K dataset. As shown in Table 13, MiLoRA outperforms LoRA and PiSSA by an average of 1.6 and 3.5 EM scores, respectively, demonstrating its effectiveness on the Qwen backbone. Additionally, we conduct experiments on LLMs such as LLaMA2 and LLaMA3 (Section 4.1), as well as vision-language models like LLaVA-1.5 (Section 4.2). Collectively, these results confirm the effectiveness of MiLoRA across diverse model architectures.
B.2 Complementary Weighet Metrics Analysis